# Spend Wisely: Maximizing Post-Training Gains in Iterative Synthetic Data Boostrapping

### v1

@@ B.1 Notations
$\mathcal{P}_{\theta}(X)\;=\;\mathcal{P}\left(f(\omega;\theta)\in X\right),\quad\text{for any measurable set }X\subseteq\mathcal{T}.$
@@
$\mathcal{P}_{\mathcal{R};\theta}(\mathcal{B})\;=\;\mathcal{P}_{\theta}\left(\mathcal{R}(x)\in\mathcal{B}\right),\quad\text{for any Borel set }\mathcal{B}\subseteq[0,1].$
@@
$r(\theta):=\int_{[0,1]}r\mathrm{d}\mathcal{P}_{\mathcal{R};\theta}(r)=\mathbb{E}_{x\sim\mathcal{P}(\theta)}\mathcal{R}(x).$
@@ B.2 Assumptions
We always consider the general case where the initial value $\theta^{(0)}$ is sufficiently accurate. The distribution $\mathcal{P}_{\theta^{(0)}}$ is assumed as non-degenerate and there exists a constant $c_{r}>0$ such that $\mathrm{d}\mathcal{P}_{r;\theta^{(0)}}\geq c_{r}$.
@@
    The covering number $\mathcal{N}(\epsilon,\Theta,\|\cdot\|_{2})$ for the parameter space $\Theta$ is bounded, which means
@@
This assumption governs both the complexity of the parameter space $\Theta$ and the smoothness of the likelihood function $l(x;\theta)$. Here, we provide two examples to show that these assumptions are reasonable. Firstly, if $\Theta$ is a compact finite-dimensional space or a functional space sufficiently smooth, then the coverage number requirement (Equation 3) is satisfied. Secondly, for a special case when $l(x;\theta)$ is the negative log-likelihood for a Gaussian distribution, the derivative condition (Equation 5) holds with $c=2$ and $\beta=1$.
@@
This assumption ensures that, within a local neighborhood of $\theta$, any parameter update that reduces the expected likelihood under $\theta_{1}$’s distribution also decreases the expected likelihood under $\theta$’s distribution, relative to $\mathbb{P}l_{\theta_{1}}(\theta_{1})$. Concretely, each iteration of the algorithm samples from $\theta_{1}$ but then updates the parameter $\theta$. Since the data remain the same while $\theta$ changes, this condition guarantees that a reduction in the loss is preserved across iterations. For instance, if $\mathcal{P}_{\theta}$ is a Gaussian distribution with mean $\theta$ and standard deviation 1, and $\mathcal{R}(x)$ is the normalized standard Gaussian density, one can verify that assumption B.2 holds with $\gamma=1/2$.
@@
This assumption establishes a direct relationship between the performance gap $(r^{*}-r(\theta))$ and the expected loss $\mathbb{P}l_{\theta}(\theta)$. Specifically, the upper bound shows that if $\mathbb{P}l_{\theta}(\theta)$ is small, then $(r^{*}-r(\theta))$ must also be small, which is crucial for proving convergence of our proposed policy. For example, in the Gaussian setting discussed in the remark of assumption B.2, one can verify that $c_{2}=1/12$. Conversely, the lower bound indicates that a small performance gap necessarily implies a correspondingly small expected loss, thereby capturing the sensitivity of the loss to improvements in the expected reward. Since both $r^{*}-r(\theta)$ and $\mathbb{P}l_{\theta}(\theta)$ quantify how well the policy performs, it is natural that each can bound the other. This relationship serves as a key underpinning for the theoretical results presented in subsequent sections.
@@ B.3 Main theorems
Under assumption B.3, there exists a constant $c>0$ such that, for any $T$ and any constant policy $\pi_{\mathrm{const}}$, with probability at least $1/4$,
@@
Under assumptions B.1, B.2 and B.3, for any increasing policy $\pi$ and $\varepsilon>0$, there exists a sufficiently large $T$ such that, with probability greater than $1-\sum_{t=0}^{T-1}n_{t}^{-4}$,
@@
Next, we demonstrate that the exponential-growth policy from Theorem B.6 achieves this convergence more effectively than polynomial growth policies, including the constant policies.
@@
Under assumptions B.1, B.2 and B.3, there exists an exponential growth policy $\pi_{\mathrm{exp}}^{*}$ such that with probability at least $1-\sum_{t=0}^{T-1}n_{t}^{-4}$,
@@
Consider the minimum number of iterations $T^{*}(\pi,\varepsilon)$ needed for a policy $\pi$ to achieve a reward that is within $\varepsilon$ of the optimal $r^{*}$ as defined in Equation 2, then the cost $C(\pi,T^{*}(\pi,\varepsilon))$ is the minimum cost to attain the target performance for policy $\pi$. The following theorem compares the total cost incurred by exponential growth policies and polynomial growth policies (including constant policies) in a worst-case scenario.
@@
Under the same conditions as in Theorem B.6, for any constant or polynomial growth policy $\pi$ and a sufficiently small $\varepsilon$, we have
@@
with probability at least $1-\sum_{t=0}^{T^{*}(\pi_{\mathrm{exp}},\varepsilon)-1}n_{t}^{-4}$. The supremum is taken over all possible parameter spaces $\Theta,\Theta^{*}$, loss functions $l(x;\theta)$, and reward models $\mathcal{R}(x)$ satisfying the same conditions as in Theorem B.6.
@@ B.4 Proofs
In this subsection, we present proofs for all statements in our theoretical results. We begin with the proof of Theorem B.6, proceed to prove Theorems B.4 and B.5, and conclude with the proof of Theorem B.7.
Proof of Theorem B.6.
@@
By assumption B.1 and a Taylor expansion, choosing the step size $\eta=1/(2\beta)$ gives
@@
$\leq-\eta\left(\nabla_{\theta}\mathbb{P}_{n_{0}}l\left(X;\theta^{(0)}\right)\right)^{\mathsf{T}}\nabla_{\theta}\mathbb{P}_{n_{0}}l\left(X;\theta^{(0)}\right)+\beta\eta^{2}\left\|\nabla_{\theta}\mathbb{P}_{n_{0}}l\left(X;\theta^{(0)}\right)\right\|^{2}_{2}$
@@
$\leq-\frac{c}{4\beta}\mathbb{P}_{n_{0}}l\left(X;\theta^{(0)}\right),$
@@
where $\varepsilon_{i}$ ($1\leq i\leq n$) are i.i.d. Rademacher random variables. By Dudley’s entropy integral and assumption B.1, we then get
[…]
@@
Combining Equations 7 and 8 with Equation 6 and choosing $t_{1}=2(N_{0}r\left(\theta^{(0)}\right))^{1/2}\log N_{0}$, we see that, with probability at least $1-N_{0}^{-4}$,
$\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(1)}\right)-\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)$
$\leq$ $-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)+\left(2-\frac{c}{4\beta}\right)\sup_{\theta\in\Theta}\left|\mathbb{P}l_{\theta^{(0)}}(\theta)-\mathbb{P}_{n_{0}}l(X;\theta)\right|$
$\leq$ $-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)+\left(2-\frac{c}{4\beta}\right)2LC_{\Theta}\log c_{\Theta}n_{0}^{-1/2}$
$\leq$ $-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)+\left(2-\frac{c}{4\beta}\right)2LC_{\Theta}\log c_{\Theta}\left(N_{0}r\left(\theta^{(0)}\right)-2\left(N_{0}r\left(\theta^{(0)}\right)\right)^{1/2}\log N_{0}\right)^{-1/2}.$
@@
Define ~$\tilde{\mathbb{P}}l_{\theta}=\mathbb{P}l_{\theta}(\theta)$, this can be rewritten as
~$\tilde{\mathbb{P}}l_{\theta^{(1)}}-\tilde{\mathbb{P}}l_{\theta^{(0)}}\leq-\frac{\gamma c}{8\beta}\tilde{\mathbb{P}}l_{\theta^{(0)}}.$
By assumption B.3,
@@
$\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t+1)}\right)-\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t)}\right)$
$\leq$ $-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t)}\right)+\left(2-\frac{c}{4\beta}\right)\sup_{\theta\in\Theta}\left|\mathbb{P}l_{\theta^{(t)}}(\theta)-\mathbb{P}_{n_{t}}l(X;\theta)\right|$
$\leq$ $-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t)}\right)+\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}n_{t}^{-1/2}$
$\leq$ $-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t)}\right)+\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}\left(N_{t}r\left(\theta^{(t)}\right)-2\left(N_{t}r\left(\theta^{(t)}\right)\right)^{1/2}\log N_{t}\right)^{-1/2}.$
@@
$\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}\left(N_{t}r\left(\theta^{(t)}\right)-2\left(N_{t}r\left(\theta^{(t)}\right)\right)^{1/2}\log N_{t}\right)^{-1/2}\leq\frac{c}{8\beta}\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)\left(1-\frac{\gamma c}{8\beta}\right)^{t}.$
@@
We choose $N_{t}$ as
$N_{t}=r\left(\theta^{(t)}\right)^{-1}\left(2\left(16\beta-2c\right)c^{-1}C_{L}C_{\Theta}\log c_{\Theta}\right)^{2}\left(\tilde{\mathbb{P}}l_{\theta}^{(0)}\right)^{-2}\left(1-\frac{\gamma c}{8\beta}\right)^{-2t}.$
For the initial value $\theta^{(0)}$ which is sufficiently accurate, we have $r\left(\theta^{(t)}\right)^{1/2}\geq r\left(\theta^{(0)}\right)^{1/2}\geq 4N_{t}^{-1/2}\log N_{t}$, ensuring that our choice of $N_{t}$ meets the needed condition.
Finally, we define $\zeta=\gamma c/(8\beta-\gamma c)$. So
@@
we obtain the results stated in Theorem B.6. $\square$
Proof of Theorem B.4. By assumption B.3, we have
$\sup_{\theta\in\Theta}\mathbb{E}_{x\sim\mathcal{P}_{\theta}}\mathcal{R}(x)-\mathbb{E}_{x\sim\mathcal{P}_{\theta_{\pi_{\mathrm{const}}}^{(T)}}}\mathcal{R}(x)$ $\geq c_{3}\mathbb{P}l_{\theta_{\pi_{\mathrm{const}}}^{(T)}}\theta_{\pi_{\mathrm{const}}}^{(T)}$
$\geq c_{3}\left(\mathbb{P}l_{\theta_{\pi_{\mathrm{const}}}^{(T)}}\theta_{\pi_{\mathrm{const}}}^{(T)}-\mathbb{P}_{n_{0}}l\left(X;\theta_{\pi_{\mathrm{const}}}^{(T)}\right)\right)+\mathbb{P}_{n_{0}}l\left(X;\theta_{\pi_{\mathrm{const}}}^{(T)}\right)$
$\geq c_{3}\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}\left(l\left(x_{i};\theta_{\pi_{\mathrm{const}}}^{(T)}\right)-\mathbb{E}_{X\sim\mathcal{P}_{l\left(x_{i};\theta_{\pi_{\mathrm{const}}}^{(T)}\right)}}l\left(X;\theta_{\pi_{\mathrm{const}}}^{(T)}\right)\right)$
$\geq c_{3}n_{0}^{-1/2}\Phi(0.75)\operatorname{Var}_{X\sim\mathcal{P}_{l\left(x_{i};\theta_{\pi_{\mathrm{const}}}^{(T)}\right)}}\left[l\left(X;\theta_{\pi_{\mathrm{const}}}^{(T)}\right)\right]^{2}$
with probability larger than $1-0.75=0.25$, where we invoke the central limit theorem for the last inequality. $\square$
Proof of Theorem B.5. Let $\varepsilon>0$ be given. We first show that there exists some $t$ satisfying
@@
Following the proof of Theorem B.6, we have
~$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t+1)}}-\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\leq-\frac{\gamma c}{4\beta}\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}+\gamma\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}n_{t}^{-1/2}.$
There exist a sufficiently large $T^{\prime}$ such that
@@
    ~$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\leq\varepsilon/(2c_{2})$.
    By assumption B.3,
@@
    ~$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}>\varepsilon/(2c_{2})$.
@@
    ~$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t+1)}}-\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\leq-\frac{\gamma c}{8\beta}\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}.$
    Repeating this argument for $t=T^{\prime},T^{\prime}+1,\cdots$ whenever ~$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}>\varepsilon/(2c_{2})$ shows that there must exist a finite $K$ (so $T=T^{\prime}+K$) where either
@@
    $\geq r^{*}-c_{2}\frac{\gamma c}{4\beta}\tilde{\mathbb{P}}l_{\theta_{\pi}^{(T^{\prime})}}\left(1-\frac{\gamma c}{8\beta}\right)^{T-T^{\prime}}$
    $\geq r^{*}-\frac{1}{2}\varepsilon,$
    or else ~$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(T)}}\leq\varepsilon/(2c_{2})$, in which case (a) implies
@@
From part (b) above, whenever ~$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\geq\varepsilon/(2c_{2})$, we have
~$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t+1)}}-\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\leq-\frac{\gamma c}{8\beta}\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}.$
@@
Proof of Theorem B.7. We prove this theorem by showing that for a given policy $\pi^{\prime}$, there exists a specified parameter space $\Theta$ and $\Theta^{*}$, a loss function $l(x;\theta)$, and a reward model $\mathcal{R}(x)$, under assumptions B.1, B.2 and B.3 such that
@@
$\geq C(\pi_{\mathrm{exp}},T^{*}(\pi_{\mathrm{exp}},\varepsilon)),$
with probability at least $1-\sum_{t=0}^{T^{*}(\pi_{\mathrm{exp}},\varepsilon)-1}n_{t}^{-4}$.
To begin, use Theorem B.6 to obtain

### v2

@@ B.1 Notations
$\mathcal{P}_{\theta}(X)\;=\;\mathcal{P}\left(f(\omega;\theta)\in X\right)\quad\text{for any measurable set }X\subseteq\mathcal{T}.$
@@
$\mathcal{P}_{\mathcal{R};\theta}(\mathcal{B})\;=\;\mathcal{P}_{\theta}\left(\mathcal{R}(x)\in\mathcal{B}\right)\quad\text{for any Borel set }\mathcal{B}\subseteq[0,1].$
@@
$r(\theta):=\int_{[0,1]}r~\mathrm{d}\mathcal{P}_{\mathcal{R};\theta}(r)=\mathbb{E}_{x\sim\mathcal{P}(\theta)}[\mathcal{R}(x)].$
@@ B.2 Assumptions
We always consider the general case where the initial value $\theta^{(0)}$ is sufficiently accurate. The distribution $\mathcal{P}_{\theta^{(0)}}$ is assumed as non-degenerate and there exists a constant $c_{r}$ such that
$\mathrm{d}\mathcal{P}_{r;\theta^{(0)}}\geq c_{r}>0.$
@@
    The covering number $\mathcal{N}(\epsilon,\Theta,\|\cdot\|_{2})$ for the parameter space $\Theta$ is bounded, which means the inequality
@@
This assumption governs both the complexity of the parameter space $\Theta$ and the smoothness of the likelihood function $l(x;\theta)$. Here, we provide two examples to show that these assumptions are reasonable. Firstly, if $\Theta$ is a compact finite-dimensional space or a functional space sufficiently smooth, then the coverage number requirement (Equation˜3) is satisfied. Secondly, for a special case when $l(x;\theta)$ is the negative log-likelihood for a Gaussian distribution, the derivative condition (Equation˜5) holds with $c=2$ and $\beta=1$.
@@
This assumption ensures that, within a local neighborhood of $\theta$, any parameter update that reduces the expected likelihood under $\theta_{1}$’s distribution also decreases the expected likelihood under $\theta$’s distribution, relative to $\mathbb{P}l_{\theta_{1}}(\theta_{1})$. Concretely, each iteration of the algorithm samples from $\theta_{1}$ but then updates the parameter $\theta$. Since the data remain the same while $\theta$ changes, this condition guarantees that a reduction in the loss is preserved across iterations. For instance, if $\mathcal{P}_{\theta}$ is a Gaussian distribution with mean $\theta$ and standard deviation 1, and $\mathcal{R}(x)$ is the normalized standard Gaussian density, one can verify that Assumption˜B.2 holds with $\gamma=1/2$.
@@
This assumption establishes a direct relationship between the performance gap $(r^{*}-r(\theta))$ and the expected loss $\mathbb{P}l_{\theta}(\theta)$. Specifically, the upper bound shows that if $\mathbb{P}l_{\theta}(\theta)$ is small, then $(r^{*}-r(\theta))$ must also be small, which is crucial for proving convergence of our proposed policy. For example, in the Gaussian setting discussed in the remark of Assumption˜B.2, one can verify that $c_{2}=1/12$. Conversely, the lower bound indicates that a small performance gap necessarily implies a correspondingly small expected loss, thereby capturing the sensitivity of the loss to improvements in the expected reward. Since both $\left(r^{*}-r(\theta)\right)$ and $\mathbb{P}l_{\theta}(\theta)$ quantify how well the policy performs, it is natural that each can bound the other. This relationship serves as a key underpinning for the theoretical results presented in subsequent sections.
@@ B.3 Main Theorems
Under Assumption˜B.3, there exists a constant $c>0$ such that, for any $T$ and any constant policy $\pi_{\mathrm{const}}$, with probability at least $1/4$,
@@
Under Assumptions˜B.1, B.2 and B.3, for any increasing policy $\pi$ and $\varepsilon>0$, there exists a sufficiently large $T$ such that, with probability greater than $\left(1-\sum_{t=0}^{T-1}n_{t}^{-4}\right)$,
@@
Next, we demonstrate that the exponential-growth policy from Theorem˜B.6 achieves this convergence more effectively than polynomial growth policies, including the constant policies.
@@
Under Assumptions˜B.1, B.2 and B.3, there exists an exponential growth policy $\pi_{\mathrm{exp}}^{*}$ such that with probability at least $\left(1-\sum_{t=0}^{T-1}n_{t}^{-4}\right)$,
@@
Consider the minimum number of iterations $T^{*}(\pi,\varepsilon)$ needed for a policy $\pi$ to achieve a reward that is within $\varepsilon$ of the optimal $r^{*}$ as defined in Equation˜2, then the cost $C(\pi,T^{*}(\pi,\varepsilon))$ is the minimum cost to attain the target performance for policy $\pi$. The following theorem compares the total cost incurred by exponential growth policies and polynomial growth policies (including constant policies) in a worst-case scenario.
@@
Under the same conditions as in Theorem˜B.6, for any constant or polynomial growth policy $\pi$ and a sufficiently small $\varepsilon$, we have
@@
with probability at least $\left(1-\sum_{t=0}^{T^{*}(\pi_{\mathrm{exp}},\varepsilon)-1}n_{t}^{-4}\right)$. The supremum is taken over all possible parameter spaces $\Theta,\Theta^{*}$, loss functions $l(x;\theta)$, and reward models $\mathcal{R}(x)$ satisfying the same conditions as in Theorem˜B.6.
@@ B.4 Proofs
In this subsection, we present proofs for all statements in our theoretical results. We begin with the proof of Theorem˜B.6, proceed to prove Theorems˜B.4 and B.5, and conclude with the proof of Theorem˜B.7.
Proof of Theorem˜B.6.
@@
By Assumption˜B.1 and a Taylor expansion, choosing the step size $\eta=1/(2\beta)$ gives
@@
$\leq-\eta\left(\nabla_{\theta}\mathbb{P}_{n_{0}}l\left(X;\theta^{(0)}\right)\right)^{\mathsf{T}}\nabla_{\theta}\mathbb{P}_{n_{0}}l\left(X;\theta^{(0)}\right)$
$\hskip 107.63855pt+\beta\eta^{2}\left\|\nabla_{\theta}\mathbb{P}_{n_{0}}l\left(X;\theta^{(0)}\right)\right\|^{2}_{2}$
@@
$\leq-\frac{c}{4\beta}\mathbb{P}_{n_{0}}l\left(X;\theta^{(0)}\right).$
@@
where $\varepsilon_{i}$ ($1\leq i\leq n$) are i.i.d. Rademacher random variables. By Dudley’s entropy integral and Assumption˜B.1, we then get
$\sup_{\theta\in\Theta}\left|\mathbb{P}l_{\theta^{(0)}}(\theta)-\mathbb{P}_{n}l(X;\theta)\right|$ $\leq 2\mathcal{R}_{\theta}$ (8)
$\leq 2Ln^{-1/2}\int_{0}^{\infty}\sqrt{\log\mathcal{N}(\epsilon,\Theta,\|\cdot\|_{2})}~\mathrm{d}\epsilon$
$\leq 2LC_{\Theta}\log c_{\Theta}n^{-1/2}.$
@@
Combining Equations˜7 and 8 with Equation˜6 and choosing
$t_{1}=2\left(N_{0}r\left(\theta^{(0)}\right)\right)^{1/2}\log N_{0},$
we see that, with probability at least $\left(1-N_{0}^{-4}\right)$,
$\begin{aligned} &\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(1)}\right)-\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)\\ \leq&-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)+\left(2-\frac{c}{4\beta}\right)\sup_{\theta\in\Theta}\left|\mathbb{P}l_{\theta^{(0)}}(\theta)-\mathbb{P}_{n_{0}}l(X;\theta)\right|\\ \leq&-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)+\left(2-\frac{c}{4\beta}\right)2LC_{\Theta}\log c_{\Theta}n_{0}^{-1/2}\\ \leq&-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)+\left(2-\frac{c}{4\beta}\right)2LC_{\Theta}\log c_{\Theta}\left(N_{0}r\left(\theta^{(0)}\right)-2\left(N_{0}r\left(\theta^{(0)}\right)\right)^{1/2}\log N_{0}\right)^{-1/2}.\end{aligned}$
@@
Define $\tilde{\mathbb{P}}l_{\theta}=\mathbb{P}l_{\theta}(\theta)$. This can be rewritten as
$\tilde{\mathbb{P}}l_{\theta^{(1)}}-\tilde{\mathbb{P}}l_{\theta^{(0)}}\leq-\frac{\gamma c}{8\beta}\tilde{\mathbb{P}}l_{\theta^{(0)}},$
and by Assumption˜B.3 we have the inequality
@@
$\begin{aligned} &\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t+1)}\right)-\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t)}\right)\\ \leq&-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t)}\right)+\left(2-\frac{c}{4\beta}\right)\sup_{\theta\in\Theta}\left|\mathbb{P}l_{\theta^{(t)}}(\theta)-\mathbb{P}_{n_{t}}l(X;\theta)\right|\\ \leq&-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t)}\right)+\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}n_{t}^{-1/2}\\ \leq&-\frac{c}{4\beta}\mathbb{P}l_{\theta^{(t)}}\left(\theta^{(t)}\right)+\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}\left(N_{t}r\left(\theta^{(t)}\right)-2\left(N_{t}r\left(\theta^{(t)}\right)\right)^{1/2}\log N_{t}\right)^{-1/2}.\end{aligned}$
@@
$\begin{aligned} \left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}\left(N_{t}r\left(\theta^{(t)}\right)-2\left(N_{t}r\left(\theta^{(t)}\right)\right)^{1/2}\log N_{t}\right)^{-1/2}\leq\frac{c}{8\beta}\mathbb{P}l_{\theta^{(0)}}\left(\theta^{(0)}\right)\left(1-\frac{\gamma c}{8\beta}\right)^{t}.\end{aligned}$
@@
Consequently, we choose
$N_{t}=r\left(\theta^{(t)}\right)^{-1}\left(2\left(16\beta-2c\right)c^{-1}C_{L}C_{\Theta}\log c_{\Theta}\right)^{2}\left(\tilde{\mathbb{P}}l_{\theta}^{(0)}\right)^{-2}\left(1-\frac{\gamma c}{8\beta}\right)^{-2t},$
then for the initial value $\theta^{(0)}$ which is sufficiently accurate, we have
$r\left(\theta^{(t)}\right)^{1/2}\geq r\left(\theta^{(0)}\right)^{1/2}\geq 4N_{t}^{-1/2}\log N_{t},$
ensuring that our choice of $N_{t}$ meets the needed condition.
Finally, we define $\zeta=\gamma c/(8\beta-\gamma c)$, then
@@
we obtain the results stated in Theorem˜B.6. $\square$
Proof of Theorem˜B.4. By Assumption˜B.3, we have
$\begin{aligned} \sup_{\theta\in\Theta}\mathbb{E}_{x\sim\mathcal{P}_{\theta}}[\mathcal{R}(x)]-\mathbb{E}_{x\sim\mathcal{P}_{\theta_{\pi_{\mathrm{const}}}^{(T)}}}[\mathcal{R}(x)]&\geq c_{3}\mathbb{P}l_{\theta_{\pi_{\mathrm{const}}}^{(T)}}\theta_{\pi_{\mathrm{const}}}^{(T)}\\ &\geq c_{3}\left(\mathbb{P}l_{\theta_{\pi_{\mathrm{const}}}^{(T)}}\theta_{\pi_{\mathrm{const}}}^{(T)}-\mathbb{P}_{n_{0}}l\left(X;\theta_{\pi_{\mathrm{const}}}^{(T)}\right)\right)+\mathbb{P}_{n_{0}}l\left(X;\theta_{\pi_{\mathrm{const}}}^{(T)}\right)\\ &\geq c_{3}\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}\left(l\left(x_{i};\theta_{\pi_{\mathrm{const}}}^{(T)}\right)-\mathbb{E}_{X\sim\mathcal{P}_{l\left(x_{i};\theta_{\pi_{\mathrm{const}}}^{(T)}\right)}}\left[l\left(X;\theta_{\pi_{\mathrm{const}}}^{(T)}\right)\right]\right)\\ &\geq c_{3}n_{0}^{-1/2}\Phi(0.75)\operatorname{Var}_{X\sim\mathcal{P}_{l\left(x_{i};\theta_{\pi_{\mathrm{const}}}^{(T)}\right)}}\left[l\left(X;\theta_{\pi_{\mathrm{const}}}^{(T)}\right)\right]^{2}\end{aligned}$
with probability larger than 10.75=0.251-0.75=0.25, where we invoke the central limit theorem for the last inequality. $\square$
Proof of Theorem˜B.5. Let $\varepsilon>0$ be given. We first show that there exists some $t$ satisfying
@@
Following the proof of Theorem˜B.6, we have
$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t+1)}}-\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\leq-\frac{\gamma c}{4\beta}\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}+\gamma\left(2-\frac{c}{4\beta}\right)2C_{L}C_{\Theta}\log c_{\Theta}n_{t}^{-1/2}.$
There exists a sufficiently large $T^{\prime}$ such that
@@
    $\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\leq\varepsilon/(2c_{2})$.
    By Assumption˜B.3, we have
@@
    $\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}>\varepsilon/(2c_{2})$.
@@
    $\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t+1)}}-\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\leq-\frac{\gamma c}{8\beta}\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}.$
    Repeating this argument for $t=T^{\prime},T^{\prime}+1,\cdots$ whenever $\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}>\varepsilon/(2c_{2})$ shows that there must exist a finite $K$ (so $T=T^{\prime}+K$) where either
@@
    $\geq r^{*}-c_{2}\frac{\gamma c}{4\beta}\tilde{\mathbb{P}}l_{\theta_{\pi}^{(T^{\prime})}}\left(1-\frac{\gamma c}{8\beta}\right)^{T-T^{\prime}}\geq r^{*}-\frac{1}{2}\varepsilon,$
    or else $\tilde{\mathbb{P}}l_{\theta_{\pi}^{(T)}}\leq\varepsilon/(2c_{2})$, in which case (a) implies
@@
From part (b) above, whenever $\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\geq\varepsilon/(2c_{2})$, we have
$\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t+1)}}-\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}\leq-\frac{\gamma c}{8\beta}\tilde{\mathbb{P}}l_{\theta_{\pi}^{(t)}}.$
@@
Proof of Theorem˜B.7. We prove this theorem by showing that for a given policy $\pi^{\prime}$, there exists a specified parameter space $\Theta$ and $\Theta^{*}$, a loss function $l(x;\theta)$, and a reward model $\mathcal{R}(x)$, under Assumptions˜B.1, B.2 and B.3 such that
@@
$\geq C(\pi_{\mathrm{exp}},T^{*}(\pi_{\mathrm{exp}},\varepsilon))$
with probability at least $\left(1-\sum_{t=0}^{T^{*}(\pi_{\mathrm{exp}},\varepsilon)-1}n_{t}^{-4}\right)$.
To begin, we use Theorem˜B.6 to obtain